OpenStudy (oliviagggg):

MEDAL FOR ANSWER PLEASE HELP Select all of the equations that will produce exponential decay. (More than one can be chosen) 1: y=(0.1)^x 2: y=2(1/4)^x 3: y=(3)^x 4: y=1/2 (5)^x 5: y=(2)^x 6: y = 3(0.4)^x

OpenStudy (mathstudent55):

In an equation of the form: \(y = ak^x\) where a is a positive number, you will have exponential decay if \(0 \le k \lt 1\)

OpenStudy (oliviagggg):

I don't get it

OpenStudy (mathstudent55):

If you raise a number larger than 1 to larger and larger exponents starting with 1, the number gets larger and larger. For example, 2^1 = 2 2^2 = 4 2^3 = 8 By raising 2 to a larger and larger exponent, the result is a larger number. Now try raising a number between 0 and 1 to larger and larger exponents. For example, raise 1/10 to larger and larger powers. (1/10)^1 = 1/10 (1/10)^2 = 1/100 (1/10^3 = 1/1000 As you can see, the numbers get smaller and smaller.

OpenStudy (mathstudent55):

Exponential decay is based on a number between 0 and 1 being raised to a larger and larger power and giving you smaller results.

OpenStudy (oliviagggg):

so i need to answer all the choices and every one that is smaller is right?

OpenStudy (mathstudent55):

Look at all the functions in your choices. They all have x as an exponent. As x becomes larger and larger, if what is being raised to x is between 0 and 1, the result will get smaller and smaller.

OpenStudy (mathstudent55):

All the choices that have a number between 0 and 1 being raised to x are exponential decay. If the number being raised to x is larger than 1, then it is exponential growth.

OpenStudy (oliviagggg):

so if x is 0 1 or in between its decay right?

OpenStudy (mathstudent55):

1: y=(0.1)^x <--- 0.1 is being raised to x 2: y=2(1/4)^x <--- 1/4 is being raised to x 3: y=(3)^x <--- 3 is being raised to x 4: y=1/2 (5)^x <--- etc. 5: y=(2)^x 6: y = 3(0.4)^x

OpenStudy (mathstudent55):

I think you're on the right track, but it's not exactly as you stated. 1. It's a number being raised to x, not x being raised to a power. 2. The number being raised to x cannot be 0 or 1. It must be in between 0 and 1 for decay.

OpenStudy (mathstudent55):

I started going through the choices above. Do you see what I did? Choices 1 and 2 have 1/10 and 1/4 raised to x. Since 1/10 and 1/4 are both between 0 and 1, choices 1 and 2 are exponential decay. Now look at choice 3. The number being raised to x is 3. 3 is not between 0 and 1, so choice 3 is not exponential decay.

OpenStudy (mathstudent55):

Now look at choices 4, 5, and 6. Which ones have a number between 0 and 1 being raised to x? The ones that do are exponential decay.

OpenStudy (oliviagggg):

are the ones in parenthesis being raised to x?

OpenStudy (oliviagggg):

because if they are the only other one i would think that is decay is 6

OpenStudy (oliviagggg):

Thank you @mathstudent55

OpenStudy (mathstudent55):

Correct. Choices 1, 2, and 6 are decay.

OpenStudy (mathstudent55):

You're welcome.